What capacitance (in pF) do you need to produce a resonant frequency of 1.50 GHz, when using a 8.00 nH inductor?
Answer #1

At resonance frequency

Inductive reactance = Capacitive reactance

XL = Xc

2*pi*f*L = 1/(2*pi*f*C)

C = 1/(4*pi^2*f0^2*L)

f0 = resonance frequency = 1.50 GHz = 1.50*10^9 Hz

L = 8 nH = 8*10^-9 H

So,

C = 1/(4*pi^2*(1.50*10^9)^2*8*10^-9)

C = 1.41*10^-12 F

C = 1.41 pF

Please Upvote.